OpenStudy (anonymous):

2 over 3x plus 1 over 3x minus 6 equals 1. what is x?

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x+\frac{ 1 }{ 3 }x-6=1 ?\]

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }+\frac{ 1 }{ 3 } = \frac{ 3 }{ 3 } \implies 1\]

OpenStudy (anonymous):

\[x-6=1 ~ \implies ~ x= 7 \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np